# Aligning LLMs for FL-free Program Repair

### v1

@@ IV-A4 Baselines
IV-B RQ1: Comparison against LLM-based APR
We compare D4C with five state-of-the-art LLM-based APR approaches on Defects4J, the most commonly used benchmark for APR. Since some of the baselines are not open-sourced, we cannot re-run them to reproduce the result. Thus, we reuse their Defects4J results reported in the original paper. Although using Defects4J may have the the risk of exposing its data to LLM’s training corpus, we still use it for evaluation because (1) existing work [17] has shown that data leakage is less of a concern for APR compared to other code-related tasks as the training corpus often contains at most the individual versions of the program (i.e., learning buggy or fixed version only), and LLMs can hardly learn the APR tasks without bug-fix pair corpus during training; and (2) the baselines have also been evaluated on Defects4J (hence, they are subject to the same data leakage risk). Following prior work, we retain this setting to ensure fairness of the comparison.
The comparison result is shown in Table II. We calculated the number of correct patches on Defects4J (v1.2, v2.0, and their sum), and the number of patches sampled per bug (notated as “Sampled”). D4C largely outperforms all the existing state-of-the-art methods provided with perfect bug locations for infilling-style APR. Specifically, D4C (GPT-4) can generate almost 10% more correct patches than ChatRepair, the latest APR methods which uses GPT-4 for multi-run dialogue with provided perfect bug locations. Additionally, D4C does not require FL tools to provide perfect bug locations, which means that D4C is a more practical tool in development, since perfect FL is difficult, if not impossible, to achieve in practice. Moreover, D4C only needs to sample 10 times for each bug in patch generation, which is 90% fewer than that required by the most efficient baseline (sampling at least 100 patches). The fewer generated patches also illustrate the efficiency of D4C in fast program repair. Therefore, the results suggest D4C is an effective and efficient APR approach. However, due to the potential of data leakage, we cannot directly verify our insight of D4C using this comparison. Thus, we conduct more in-depth experiments on DebugBench in Sec. IV-C and Sec. IV-D.
@@ IV-C RQ2: Effectiveness of the two insights
RQ2 aims to validate the effectiveness of the two insights introduced in Sec. II, and how the two insights contribute to the overall performance improvement in D4C. We design two additional questions for the insights: (IQ1) Does changing output to complete function align the training objectives and result in better performance? (IQ2) Does providing artifacts to the model enable an end-to-end APR and achieves a better performance? Specifically, we evaluate the two insights on DebugBench since there is no threat of data leakage. Thus, we use verified patches (i.e., patches verified by LeetCode unseen tests) in DebugBench for the evaluation.
@@ IV-C1 Validation of Insight 1
To answer the IQ1 question, we calculate the average perplexity (i.e., measures how well a model predicts samples) for various output format on the white-box model (i.e., Mixtral-MoE). Lower perplexity values indicate the model is more confident at predicting a given sequence. Specifically, we measure perplexity by calculating the loss of CLM objective (Eq. 2). Measuring perplexity is the most straightforward approach to evaluate whether the generated content aligns with the distribution of training corpus, as the LLM is trained to minimize the CLM objective. Specifically, we adopt two output formats, including the fixed hunk patches that is used by infilling-style APR (denoted as Hunk) and the full-function output used by D4C (Func). Fig. 6 shows the examples of these output format. In our implementation, we also carefully modified the system instructions to adapt to each output setting. To control variables, we also calculate the perplexity of generated output (O), and the entire input&output (IO). Thus, we can compare the perplexity among different output format to validate our insight, ignoring the difference between each input format.
TABLE III: Comparison of different settings of D4C on DebugBench. We measure perplexity and verified patches to validate insights.
@@
Table III shows the results. We observe that when the output is an entire function, the output perplexity is much lower than using the fixed hunks as the output, regardless of the input. This directly validates that generating a complete function is the better way to align the training objective. To further validate that it is the objective alignment that helps D4C achieve a better APR performance, we evaluate the verified patches generated in each case. The result illustrates that using complete function as output can result in generation of more verified patches with lower perplexity than using discrete hunks. Moreover, when only generating the fixed hunks as the output, the output perplexity is much higher than its perplexity calculated with the whole input and output sequence. This indicates that completing discrete hunks only is misaligned with LLM’s training objective. Overall, these findings illustrate the effectiveness of our first insight.
@@ IV-C2 Validation of Insight 2
To answer the IQ2 question, we evaluate the patch generation ability on DebugBench using two different input formats, i.e., the bug report (notated as Report) and the buggy program where the buggy hunks are masked (Mask). To control variables, we also conducted experiments using the two previously established output formats as shown in Fig. 6. For the generated patches by Report-Hunk, as we do not provide a specific location to modify and cannot precisely determine where the generated patch should be applied, we instruct Mixtral-MoE to help us to automatically apply the generated patches to the original buggy program before patch validation. Table III shows the number of generated verified patches on both Mixtral-MoE and GPT-4.
The results indicate that Report-Func is the most effective among all evaluated groups, including Mask-Hunk, which is the input-output format of infilling-style APR. This phenomenon aligns with prior evaluations on Defects4J (RQ1). It is worthwhile to note in the experiment in Sec. IV-B, the effectiveness of D4C stems not only from the bug report design but also the objective alignment of the output. However, even when we use the same output format in this experiment, the performance of using bug report as input still outperforms that of using program where the buggy locations are masked. This suggests that bug report helps not only in fault localization but also give more information for patch generation. Thus, this phenomenon validates our second insight.
@@ IV-D RQ3: Ablation study and sensitive analysis
TABLE IV: The number of patches generated for our ablation study and sensitive analysis. We measure plausible patches for Defect4J and patches verified via unseen tests for DebugBench.
@@
Table IV shows the results of our ablation study and sensitive analysis. Overall, the results show that each component is important in guiding D4C in generating plausible/verified patches across the two benchmarks. We also observe that the “- Mask” baseline are the least effective because D4C cannot use any information from the artifact to fix the defect. Meanwhile, among the first three baselines where we remove one information from the prompt, we notice that program documents (w/o Document) are the least effective as D4C is able to generate the greatest number of plausible/verified patches without using the document. This is expected as the error message and test are more closely related to the defect than the comments that merely describe general code features.
@@ V-A Threats to Validity
External Threats. During the manual patch correctness analysis, we follow the procedure of previous work [14, 24, 19, 63]. To mitigate the potential bias in the manual analysis, two authors independently confirm the patch correctness. Any patches with disagreement were presented to the third author for review. However, since we could not submit the patch to the developers of the corresponding repository for confirmation, this threat could not be completely eliminated. We also release our results for public verification. As the effectiveness of the model may vary in different settings, our results may not generalize beyond the studied settings and other programming languages beyond the supported ones in Defect4J and DebugBench. We mitigate this threat by reusing configurations in prior work, and widely used benchmarks.
Internal Threats. One threat to internal validity lies in the correctness of the implementation and scripts that can affect our reported results. To mitigate this threat, we have carefully reviewed all the source code and scripts. Another internal threat comes from the potential data leakage problem (i.e., using Defects4j in RQ1, which may have been incorporated in LLM pre-training). We have discussed this threat in RQ1 (i.e., data leakage is not significant on APR [17], and the baselines have also been evaluated on Defects4J). To eliminate this threat in our insight validation, we selected DebugBench, a latest, leakage-free benchmark, for evaluation. This not only keeps a fair comparison in RQ1, but also ensures the credibility of our conclusions in RQ2 and RQ3.
Figure 7: A plausible patch generated at developer patch locations (Jsoup-19).
@@ V-B Qualitative Analysis of Plausible Patches
Figure 8: A correct patch generated at non-developer patch locations (JxPath-8).
@@
Table IV shows that D4C generates a total of 210 plausible patches but 30 are incorrect with respect to the human patches. Fig. 7 shows one example of the incorrect patch. Since the failed test contains the "src" tag as input, LLM can naturally generate patches that only pass the test containing "src" without considering patch robustness. This scenario is similar to the overfitting problem in APR [65]. To check for the overfitting problem, two of the authors independently evaluate the patch correctness, followed by a third person to review and confirm their disagreement to ensure the accuracy of the validation. In our manual validation, we also found that when D4C can generate five or more plausible patches for a buggy program in 10 samplings, most of these patches are correct. This observation aligns with previous work (i.e., responses generated multiple times tend to be more reliable [66]). DebugBench does not suffer from the overfitting problem because the given test examples are not used in its unseen test suite when checking for patch correctness.
@@
As we allow D4C to debug (instead of relying on fault localization), we also observe that some correct patches have been applied at different locations compared to the developer patches. Fig. 8 shows an example where the developer patch modifies two if-conditions so it is a 2-hunk bug in the dataset. However, D4C only needs to insert one line, which combines the two conditions and results in a correct patch that is semantically equivalent to the developer patch. Furthermore, if we ask GPT-4 to generate patches at these two hunks, GPT-4 can only write two Exception statements, and fails to pass the unit tests. This example shows that allowing LLMs to generate patches without restricting the location may provide more flexibility, allowing generation of correct patches.
@@ VI Conclusion
This paper presents a new approach to adapt LLMs to automated program repair. Our key insight is that the effectiveness of LLM-based APR can be enhanced by (1) aligning the output to their training objective and (2) allowing them to refine the whole program without first performing fault localization. Based on this insight, we designed D4C, an LLM-based prompting framework for APR. Our evaluation shows that D4C outperforms the SOTA APR methods with perfect fault localization by 10%. Our findings call for adopting a new paradigm and debugging workflow for future LLM-based APR approaches.

### v2

@@ IV-B RQ1: Comparison against LLM-based APR
TABLE III: The number of correct patches generated in Perfect FL and Statistical FL on Defects4J v1.2.
Method Model Perf. Stat. Drop Sampled
AlphaRepair CodeBERT 52 36 30.8% 5000
RAP-Gen CodeT5 72 48 33.3% 100
D4C GPT-4 84 80 4.8% 10
In Table II, we calculated the number of correct patches on Defects4J (v1.2, v2.0, and their sum) and patches sampled per bug (notated as “Sampled”) in the perfect localization settings. Specifically, for baselines, the perfect buggy hunks within the buggy function is provided, while for D4C, only the buggy function is provided. Notably, D4C largely outperforms all the existing state-of-the-art methods provided with perfect buggy hunks for infilling-style APR. D4C (GPT-4) can generate almost 10% more correct patches than ChatRepair, the latest APR methods which uses GPT-4 for multi-run dialogue with provided buggy hunks. Moreover, D4C only needs to sample 10 times for each bug in patch generation, which is 90% fewer than that required by the most efficient baseline (sampling at least 100 patches). The fewer generated patches also illustrate the better efficiency of D4C and less waste of cost in patch validation. However, using only perfect FL settings does not demonstrate the flexibility and practicality of D4C, which does not rely on given buggy hunks. To illustrate its effectiveness in real-world scenarios, we further conducted experiments of D4C and baselines using existing statistical FL tools.
In Table III, we report the number of correct patches on Defects4J under perfect FL (Perf.) and statistical FL (Stat.) settings. The baselines use statement-level FL tools (e.g., Tarantula [29]) to identify buggy hunks, whereas D4C employs a method-level FL tool (i.e., FLUCCS [58]) to locate buggy functions. We compare D4C only with AlphaRepair and RAP-Gen on Defects4J v1.2 for two reasons: (1) these are the only methods that reported the evaluation results with non-perfect FL setting on Defects4J v1.2 in the paper, and (2) FLUCCS’s code supports only Defects4J v1.2. In our experiments, D4C generates only 10 patches for the most suspicious faulty function (Top@1 method) given by FLUCCS. Notably, D4C repairs only 4 fewer bugs compared to the perfect FL settings, with a minimal drop of 4.8% in repair success. This is significantly lower than the infilling-style APR baselines, which experienced at least a 30% drop. These results suggest that D4C is less affected by inaccuracies in FL since locating buggy functions is easier than locating buggy hunks.
Overall, the results in Table II and Table III suggest that D4C is effective and efficient. However, due to the potential of data leakage, we cannot directly verify our insight of D4C via comparing current approaches. Thus, we conduct more in-depth experiments on DebugBench in Sec. IV-C and Sec. IV-D.
@@ IV-C RQ2: Effectiveness of the two insights
RQ2 aims to validate the effectiveness of the two insights introduced in Sec. II, and how the two insights contribute to the overall performance improvement in D4C. We design two additional questions for the insights: (IQ1) Does allowing LLM to generate a complete refined function better align its training objectives and result in better performance? (IQ2) Does providing the artifacts enable LLM to locate and repair buggy hunks itself and achieves a better performance? Specifically, we evaluate the two insights on DebugBench since there is no threat of data leakage. Thus, we use verified patches (i.e., patches verified by LeetCode unseen tests) in DebugBench for the evaluation.
@@ IV-C1 Validation of Insight 1
To answer the IQ1 question, we calculate the average perplexity, the inference loss of CLM objective (Eq. 2), for various output format on the white-box model (i.e., Mixtral-MoE). Measuring perplexity is the most straightforward approach to evaluate whether the generated content aligns with the distribution of training corpus, as LLMs are trained to minimize the CLM objective [13, 41]. Thus, a lower perplexity value indicate the model is more confident at predicting a given sequence. Specifically, we adopt two output formats, including the fixed hunk patches that is used by infilling-style APR (denoted as Hunk) and the full-function output used by D4C (Func). Fig. 6 shows the examples of these output format. In our implementation, we also carefully modified the system instructions to adapt to each output setting. To control variables, we also calculate the perplexity of generated output (O), and the entire input&output (IO). Thus, we can compare the perplexity among different output format to validate our insight, ignoring the difference between each input format.
TABLE IV: Comparison of different settings of D4C on DebugBench. We measure perplexity and verified patches to validate insights.
@@
Table IV shows the results. We observe that when the output is an entire function, the output perplexity is much lower than using the fixed hunks as the output, regardless of the input. This directly validates that generating a complete function is the better way to align the training objective. To further validate that it is the objective alignment that helps D4C achieve a better APR performance, we evaluate the verified patches generated in each case. The result illustrates that using complete function as output can result in generation of more verified patches with lower perplexity than using discrete hunks. Moreover, when only generating the fixed hunks as the output, the output perplexity is much higher than its perplexity calculated with the whole input and output sequence. This indicates that completing discrete hunks only is misaligned with LLM’s training objective. Overall, these findings illustrate the effectiveness of our first insight.
@@ IV-C2 Validation of Insight 2
To answer the IQ2 question, we evaluate the patch generation ability on DebugBench using two different input formats, i.e., the bug report (notated as Report) and the buggy program where the buggy hunks are masked (Mask). To control variables, we also conducted experiments using the two previously established output formats as shown in Fig. 6. For the generated patches by Report-Hunk, as we do not provide a specific location to modify and cannot precisely determine where the generated patch should be applied, we instruct Mixtral-MoE to help us to automatically apply the generated patches to the original buggy program before patch validation. Table IV shows the number of generated verified patches on both Mixtral-MoE and GPT-4. The results indicate that Report-Func is the most effective among all evaluated groups, including Mask-Hunk, which is the input-output format of infilling-style APR. This phenomenon aligns with prior evaluations on Defects4J (RQ1). It is worthwhile to note in the experiment in Sec. IV-B, the effectiveness of D4C stems not only from the bug report design but also the objective alignment of the output. However, even when we use the same output format in this experiment, the performance of using bug report as input still outperforms that of using program where the buggy hunks are masked. This suggests that bug report helps not only in locating the buggy hunks, but also give more information for patch generation. Thus, this phenomenon validates our second insight.
@@ IV-D RQ3: Ablation study and sensitive analysis
TABLE V: The number of patches generated for our ablation study and sensitive analysis. We measure plausible patches for Defects4J and patches verified via unseen tests for DebugBench.
@@
- Pure: D4C without including any guidance like mask or artifacts to identify bugs in the prompt, i.e., an aligned LLM.
Table V shows the results. Overall, the results show that each component is important in guiding D4C in generating plausible/verified patches across the two benchmarks. We also observe that the “- Mask” baseline are the least effective among all baselines with bug location guidance because D4C cannot use any information from the artifact to fix the defect. Meanwhile, among the first three baselines where we remove one information from the prompt, we notice that program documents (w/o Document) are the least effective as D4C is able to generate the greatest number of plausible/verified patches without using the document. This is expected as the error message and test are more closely related to the defect than the comments that merely describe general code features. Without guidance on buggy locations, pure GPT-4 can still fix bugs by following the aligned one-shot example pair to generate a refined function. While its effectiveness on Defects4J might be due to potential data leakage during pre-training, DebugBench is not affected by this issue. Despite this, GPT-4 with the completion objective still fixes more bugs (414 vs 408) on DebugBench than GPT-4 with the infilling objective and masked hunk (“Mask-Hunk” in Table IV). This suggests LLMs have learned to write bug-free code during pre-training. By aligning the inference task with the pre-trained code completion task, we can effectively leverage their pre-trained knowledge to produce high-quality code.
@@ V-A Threats to Validity
External Threats. One external threat comes from the potential data leakage problem (i.e., using Defects4J in RQ1, which may have been incorporated in LLM pre-training). We have discussed this threat in RQ1 (i.e., data leakage is not significant on APR [17], and the baselines have also been evaluated on Defects4J). To eliminate this threat in our insight validation, we selected DebugBench, a latest, leakage-free benchmark, for evaluation. This not only keeps a fair comparison in RQ1, but also ensures the credibility of our conclusions in RQ2 and RQ3. As the effectiveness of the model may vary in different settings, our results may not generalize beyond the studied settings and other programming languages beyond the supported ones in Defects4J and DebugBench. We mitigate this threat by reusing configurations in prior work, and widely used benchmarks.
Internal Threats. An internal threat lies in the incomplete test coverage problem [67] of Defects4J, which do not guarantee the correctness of test-passing patches. To mitigate this threat, we follow the previous work [14, 25, 21, 62] to manually check the correctness of plausible patches. To mitigate the potential bias in the manual analysis, two authors independently confirm the patch correctness. Any patches with disagreement were presented to the third author for review. However, since we could not submit the patch to the developers of the corresponding repository for confirmation, this threat could not be completely eliminated. Therefore, we decide to release our experimental results for public verification.
Figure 7: A plausible patch generated at developer patch locations (Jsoup-19)
@@ V-B Qualitative Analysis of Plausible Patches
Figure 8: A correct patch generated at non-developer patch locations (JxPath-8)
@@
Table V shows that D4C generates a total of 210 plausible patches but 30 are incorrect with respect to the human patches. Fig. 7 shows one example of the incorrect patch. Since the failed test contains the "src" tag as input, LLM can naturally generate patches that only pass the test containing "src" without considering patch robustness. This scenario is similar to the overfitting problem in APR [69]. To check for the overfitting problem, two of the authors independently evaluate the patch correctness, followed by a third person to review and confirm their disagreement to ensure the accuracy of the validation. In our manual validation, we also found that when D4C can generate five or more plausible patches for a buggy program in 10 samplings, most of these patches are correct. This observation aligns with previous work (i.e., responses generated multiple times tend to be more reliable [70]). DebugBench does not suffer from the overfitting problem because the given test examples are not used in its unseen test suite when checking for patch correctness.
@@
As we allow D4C to debug without given buggy hunks, we also observe that some correct patches have been applied at different locations compared to the developer patches. Fig. 8 shows an example where the developer patch modifies two if-conditions so it is a 2-hunk bug in the dataset. However, D4C only needs to insert one line, which combines the two conditions and results in a correct patch that is semantically equivalent to the developer patch. Furthermore, if we ask GPT-4 to generate patches at these two hunks, GPT-4 can only write two Exception statements, and fails to pass the unit tests. This example shows that allowing LLMs to generate patches without restricting the location may provide more flexibility, allowing generation of correct patches.
V-C Token Cost and Price of Patch Generation
The cost of patch generation is a vital criteria to evaluate D4C’s practicality. For Defects4J, the average input prompt length is 1,386.84 tokens overall, with 1,325.68 for resolved bugs and 1,369.72 for unresolved bugs. The average output completion length is 314.39 tokens, with 285.98 for resolved bugs and 325.80 for unresolved bugs. Given OpenAI’s pricing of \$0.01 per 1,000 input tokens and \$0.03 per 1,000 output tokens [68], the average cost per patch is \$0.023 (\$0.021 for resolved bugs and \$0.024 for unresolved bugs). Generating 10 patches per bug results in a total cost of \$0.23 for the entire dataset. Notably, D4C fixed 180 bugs in Defects4J, with the correct patch being the 4.29th out of 10 patches generated per bug on average (std.=1.74). Thus, if we stop patch generation once a correct patch is found, the average cost per bug will be reduced to \$0.18.
V-D Wasted Effort of Patch Validation
The wasted effort of patch validation is another criteria to evaluate D4C’s practicality. Due to weak test oracles, not all test-passing (plausible) patches are correct [71], and may have anti-patterns [67]. Thus, there is a labor cost of manually validating the correctness of plausible patches beyond the token cost of patch generation. Specifically, D4C generates an average of 5.80 plausible patches per bug (std.=3.28), with every 2.06th plausible patch being correct (std.=1.49). Compared to infilling APR methods which often require users to manually verify over 400 plausible patches before finding a correct patch with patch ranking (and over 600 without ranking) [14], D4C is significantly more user-friendly. Even in the theoretical worst case, users only need to validate up to 10 patches per bug, which is significantly fewer than with existing methods. As discussed in Sec. III-E, this efficiency explains why D4C does not require the additional re-ranking mechanisms to facilitate patch validation. Moreover, there is also a time cost for running unit tests during automated patch validation. However, D4C requires running tests only up to 10 minutes per bug, with a timeout of 1min per patch. This is much fewer than a search time limit of up to 5 hours used in previous work [14, 25]. This illustrates D4C’s efficiency.
11footnotetext: Data Availability: Our source code and experimental results are publicly available at https://github.com/CUHK-Shenzhen-SE/D4C
@@ VI Conclusion
This paper presents a new approach to adapt LLMs to automated program repair. Our key insight is that the effectiveness of LLM-based APR can be enhanced by (1) aligning the output to their training objective and (2) allowing them to refine the whole program without given buggy hunks. Based on this insight, we designed D4C, an LLM-based prompting framework for APR. Our evaluation shows that D4C outperforms the SOTA APR methods with perfect FL by 10%. Our findings call for adopting a new paradigm and debugging workflow for future LLM-based APR approaches.
VII Acknowledgment
This paper was supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2024A1515010145) and the Shenzhen Science and Technology Program (No. ZDSYS20230626091302006)